# OriStitch: A Machine Embroidery Workflow to Turn Existing Fabrics into Self-Folding 3D Textiles

### v1

@@ Abstract.
OriStitch is a computational fabrication workflow to turn existing flat fabrics into self-folding 3D structures. Users turn fabrics into self-folding sheets by machine embroidering functional threads in specific patterns on fabrics, and then apply heat to deform the structure into a target 3D structure. OriStitch is compatible with a range of existing materials (e.g., leather, woven fabric, and denim).
We present the design of specific embroidered hinges that fully close under exposure to heat. We discuss the stitch pattern design, thread and fabric selection, and heating conditions. To allow users to create 3D textiles using our hinges, we create a tool to convert 3D meshes to 2D stitch patterns automatically, as well as an end-to-end fabrication and actuation workflow. To validate this workflow, we designed and fabricated a cap (303 hinges), a handbag (338 hinges), and a cover for an organically shaped vase (140 hinges).
In technical evaluation, we found that our tool successfully converted 23/28 models (textures and volumetric objects) found in related papers. We also demonstrate the folding performance across different materials (suede leather, cork, Neoprene, and felt).
@@ 1. Introduction
Folding allows fabricating 3D structures from 2D sheets, this offers a range of advantages in fabrication workflows such as speeding up fabrication (An et al., 2018), reducing material consumption (Narumi et al., 2023), flat packing objects for efficient transportation (Tao et al., 2021), embedding sensors (Wang et al., 2020) and actuators (Yao et al., 2015) in otherwise passive structures. For the reasons above, as well as aesthetic effects, folding is a popular technique used when fabricating textiles.
@@
A common form of industrial folding is machine pleating (Zaki et al., 2021), while efficient and cheap, it only supports straight so called ”simple folds”. More advanced folding enables rich textures such as the popular smocking pattern or textiles that maintain a 3D shape. These forms of folding require more elaborate fabrication, such as weaving (Lind, 2019) or embroidery (Cane and Hodges, 1985; Gostelow, 2003; Hall, 2008; Luo et al., 2023). The resulting textures and shapes are interesting, but involve laborious manual processes.
@@
To achieve accessible workflows for folding of textiles, research focused on additive textile manufacturing (Keefe et al., 2022). This includes techniques where hinge patterns are added to fabrics either by printing (Koch et al., 2021), or embroidery (Stoychev et al., 2017). These hinges, in turn, fold using shape memory alloys (Nabil et al., 2019), actuated fibers (Nabila et al., 2023a), or even by controlling the fabrication of the fiber itself (Forman et al., 2023). While these techniques are clever, they are hard to practically implement as it requires manually laying out many individual hinges in machine instructions. Inkjet 4D Print (Narumi et al., 2023) is an automated, high-resolution folding process. The technique locally coats heat-shrinkable materials with ink to control folding angles. While the resulting geometry looks great, the resulting hinges unfold easily, and it only works on those shrinkable materials (and works best with plastic), limiting practical applicability in textile manufacturing.
@@
We propose OriStitch, a partially automated workflow to turn existing fabrics into self-folding structures through embroidery. Inherent to machine embroidery and the use of heat shrinking threads, this method is compatible with most existing off-the-shelf textiles, unlike weaving or knitting where a new textile is created from scratch, or relying on textiles that exhibit rare properties. The hinge tucks away material between faces when heat-actuated through the use of a shrinking thread, similar to 4D Embroidery (Nabila et al., 2023a). The specific design of OriStitch hinges allows them to fully close and lock in place. By laying out hinges on a 2D sheet and closing them, OriStitch allows converting flat-fabricated sheets into 3D structure. To support the laying out of hinges, we make an extension to Origamizer (Tachi, 2010) to convert 3D meshes into fabrication-ready stitch patterns. The resulting workflow allows creating 3D structures such as the cap shown in Figure 1 as well as other custom textiles like the vase cover and handbag shown in Figure 18. OriStitch integrates well with other textile manipulation techniques, such as quilting, embroidery, or sewing, allowing it to fit in with existing production workflows without reducing the expressiveness of design.
@@
Through technical validation, we find that OriStitch hinges fabricate and fold reliably using a range of materials (suede leather, cork, Neoprene, and felt). Our proposed software tool successfully converted 23/28 models found in related papers (Guseinov et al., 2017; Narumi et al., 2023; Panetta et al., 2021; Jourdan et al., 2023) to machine instructions. With this contribution, we take a small step to make the workflow of machine embroidery self-folding structures more accessible. The improved software support may allow fashion designers and discretionary users who are not mathematically inclined to design more elaborate patterns.
@@ 2. Contributions, Benefits, and Limitations
This paper makes three contributions. (1) We present a workflow, OriStitch, to create self-folding 3D geometry from flat fabrics. (2) We present a design of stitch patterns to achieve the fully closing hinges under heat treatment. And finally, (3) we present a software tool which extends Origamizer (Tachi, 2010) to convert 3D meshes to fabrication instructions. Combined, these contributions allow users to create a wide variety of shapes in a mostly automatic workflow.
Inherent to the use of developable surfaces, this process is limited to 3D objects that are topologically disks. We demonstrate that with cutting of the 3D shape and stitching together separate sheets, more complex topology (such as topological spheres and tori) are attainable. Furthermore, to be able to place hinges, objects are suspect to a resolution of edge lengths of minimal 8.4mm. Our workflow uses high temperatures (¿350∘F), as well was exposure to water, therefore the used fabric has to be robust to these conditions.
@@ 3.1. Textile Folding
Manual Processes. Manual techniques, such as hand pleating allow folding a range of materials. During hand pleating, fabric is placed between two halves of a cardboard mold, folded, and afterwards it is exposed to high temperatures. Another technique is hand-sewing continuous threads into textiles and pulling them together. Despite providing greater design freedom, these processes are labor-intensive and time-consuming, limiting their practicality.
@@ 3.3. Techniques for Self-Folding
A wide variety of fabrication processes have recently been explored to achieve complex 3D shapes through 2D configurations. One of the most common self-folding techniques is to print thermoplastic composites using Fused Deposition Modeling (FDM) 3D Printers, also known as 4D printing (An et al., 2018). Within textiles, similar methods create self folding structures by printing on pre-stretched sheets in a process called textile additive manufacturing (Brenken et al., 2018). Keefe et al. (2022) present a great overview of these techniques highlighting embroidery as one such process.
@@ 4.1. Stitch pattern design
Figure 3 shows the elementary hinge and its actuation principle. The hinges fold along the valley fold on the reflecting axis, so that the space between faces is completely tucked behind surface. As a result, the decomposed faces will be brought together to form the target 3D geometry. To accomplish this, the active thread is made of a special polyester (PE) thread (Chizimi)222https://thesewingplace.com/chizimi-shrinking-thread-300m which shrinks by 30% of its original length when exposed to heat (350∘F). The two ends of the heat shrinking thread are fixed on a pair of adjacent faces by so called lock stitches. So that under heat treatment, as the active thread shrinks, the two faces will join. Each edge contains two parallel shrinking threads to uniformly close hinges. The so called channel stitches guide the shrinking thread on the faces. The fold stitches are zigzag patterns with non-shrinking thread on top and shrinking threads on bottom (aka bobbin) for mountain folds, and vice-versa for valley folds. As the thread shrinks, the edge pre-creases in the predefined orientation, making it easier for the shrinking thread to close the hinge. Finally, the orange, water-soluble thread holds the active threads in place during fabrication.
@@ 4.1.1. Variation of hinge types
The hinge design is limited in space between two faces, as the active thread shrinks by up to 30%. Widening the hinge means more of the active thread has to be stitched to each face. To accommodate for this, OriStitch has four different hinge types displayed in Figure 4, accommodating for increasingly wide hinges by (type 2) wrapping the shrinking thread around both faces to share lock stitches (type 3) wrapping two paths of shrinking thread next to each other all the way around each of the faces and (type 4) inserting an additional pair of hinges to produce a harmonica fold and a shrinking thread that spans all the way around the other hinge.
@@
As illustrated in Figure 5, each face has up to three neighboring faces, so they contain the shrinking patterns for up to three hinges. The fold stitches bend along the edges. This avoids intersections between the active threads, allowing them to share channel stitches and lock stitches. Besides simplifying the layout, sharing channel and lock stitches speeds up fabrication.
@@ 4.1.2. Needle points
OriStitch’ hinge design requires careful placement of needle points, the stitches where bobbin threads (from below) connect with the threads on top. As illustrated in Figure 6a, in conventional embroidery, designers prioritize the top side of the fabric for aesthetic reasons, they provide extra tension on the bobbin thread, pulling it through the sheet. OriStitch cannot afford that because the active threads need to smoothly adjust their length. OriStitch adds additional needle points to distribute the tension across the thread, as shown in Figure 6b. This is specifically important at corners, where the thread on top makes a turn and would tear out the bobbin threads. OriStitch adds additional needle points at all corners and has less on the straight paths, ensuring both reliable fabrication and actuation.
@@ 4.1.3. Resolution
OriStitch is subject to a minimum size for each face to allow sufficient space to place all stitches. The minimal size of faces to be converted to OriStitch hinges has an edge length of 8.4mm and is predominantly constrained by embroidery machine limits. Typical machines have a needle point safety margin (distance between points) of 0.3mm, as shown in Figure 7, this is the main constraint for the size of faces in OriStitch. The channel stitch in the center is reduced to the minimal size, each corner requires three needle points, so on both sides of the active thread path there are exactly three points. The active thread path contains up to 6 parallel threads, which in order for the threads to stay flexible requires 0.6mm of width. The final constraining factor is the locks titch design, which again for safety margin reasons cannot be further reduced. The main conclusion is that the resolution really is machine-specific when it comes to the safety margins. That said, this assumes a dense fabric (leather, felt, or materials with tightly woven structures), if the fabric has a structure with wider gaps (typical for loosely woven or knitted structures), we recommend increasing the minimum edge length in the model past this resolution before converting to stitch patterns.
@@ 4.2.1. Materials
The core material we are using is a special polyester (PE) thread (Chizimi) that shrinks around 30% under heat treatment333https://thesewingplace.com/chizimi-shrinking-thread-300m/. We use water soluble material444https://www.fjx.co.jp/english/catalog/household/water.html to temporarily hold other threads in place before actuation. Most of our testing we have done with ”Aida cloth, 18 count” fabric555https://www.michaels.com/product/10357725, but we find in the technical evaluation that the hinges continue to work with a range of fabrics. That said, caution should be placed on the selection of fabrics. During the actuation process both exposure to wetness and high temperatures limits the range of effective materials. We also found in our evaluation that the materials need to have some stiffness to work reliably, otherwise the hinges fail to fold as expected. OriStitch comes with a calibration patch shown in Figure 10, which we recommend to use to test the required material properties as well as actuation temperature.
@@
Next to machine-embroidery, OriStitch’ workflow makes use of lasercutting. As most fabrics are relatively flexible, they require scoring or pre-creasing of the edges of the well-defined folds OriStitch expects. To reduce the force required to close the hinges and to create well-defined edges, mountain and valley folds are designed to be cut along most of the edges, with only the top, middle, and bottom sections connected as demonstrated in Figure 4, The excess material between vertices of the mesh is also cut away to make it easier to fold. While the cutting step is not strictly required for the pattern to fold (as the original layout as generated by Origamizer (Tachi, 2010) folds fine with paper without cutting), it simplifies the 2D layout as there is no need for additional patterns to fold the material away (aka ”vertex tucking molecules”), and it reduces the forces required to close each hinge.
As shown in Figure 8, OriStitch adds little tabs in the cutout areas to keep the sheet together, this minimizes the risk of the fabric folding on top of itself during embroidery and it keeps the tension in the sheet more even. Within the holes, OriStitch cuts small squares which are used for registration in the multi-hooping process. (b) OriStitch embroiders the same squares on the stabilizer sheet before attaching the fabric to the hoop. When attaching the fabric these rectangles help align both sheets. Alternatively to laser cutting, some embroidery machines use cutting needles which can be used to remove the excess material. OriStitch exports both the laser cutting plan and stitch patterns as vector graphics (SVG).

### v2

@@ Abstract.
OriStitch is a computational fabrication workflow to turn existing flat fabrics into self-folding 3D structures. It uses machine embroidering of functional threads in specific patterns on fabrics, users then apply heat to deform the structure into a target 3D structure. OriStitch is compatible with a range of existing materials (e.g., leather, woven fabric, and denim).
We present the design of specific embroidered hinges that fully close under exposure to heat. We discuss the stitch pattern design, thread and fabric selection, and heating conditions. To allow users to create 3D textiles using our hinges, we create a tool to convert 3D meshes to 2D stitch patterns automatically, as well as an end-to-end fabrication and actuation workflow. To validate this workflow, we designed and fabricated a range of application models consisting of up to 338 hinges each (the cap of Figure 1, a handbag, and a vase cover).
In technical evaluation, we found that our tool successfully 26/28 models found in related papers. We also demonstrate the folding performance across different materials (suede leather, cork, Neoprene, and felt).
@@ 1. Introduction
Folding allows fabricating 3D structures from 2D sheets, which offers a range of advantages in fabrication workflows such as speeding up fabrication (An et al., 2018), reducing material consumption (Narumi et al., 2023), flat packing objects for efficient transportation (Tao et al., 2021), embedding sensors (Wang et al., 2020) and actuators (Yao et al., 2015) in otherwise passive structures. For the reasons above, as well as aesthetic effects, folding is a popular technique used when fabricating textiles.
@@
A common form of industrial folding is machine pleating (Zaki et al., 2021), while efficient and cheap, it only supports straight so-called ”simple folds”. More advanced folding enables rich textures such as the popular smocking pattern or textiles that maintain a 3D shape. These forms of folding require more elaborate fabrication, such as weaving (Lind, 2019) or embroidery (Cane and Hodges, 1985; Gostelow, 2003; Hall, 2008; Luo et al., 2023). The resulting textures and shapes are interesting, but involve laborious manual processes.
@@
To achieve accessible workflows for folding of textiles, research focused on additive textile manufacturing (Keefe et al., 2022). This includes techniques where folding patterns are added to fabrics either by printing (Koch et al., 2021), or embroidery (Stoychev et al., 2017). The folds are actuated using shape memory alloys (Nabil et al., 2019), actuated fibers (Nabila et al., 2023a), or even by controlling the fabrication of the fiber itself (Forman et al., 2023). While these techniques are clever, they are hard to practically implement as they require manually laying out many individual hinges in machine instructions. Inkjet 4D Print (Narumi et al., 2023) is an automated, high-resolution folding process. The technique locally coats heat-shrinkable materials with ink to control folding angles. While the resulting geometry looks great, the resulting hinges unfold easily, and it only works on those shrinkable materials (and works best with plastic), limiting practical applicability in textile manufacturing.
@@
We propose OriStitch, an end-to-end workflow to turn existing fabrics into self-folding structures through embroidery. Inherent to machine embroidery, this method is compatible with most existing off-the-shelf textiles, unlike weaving or knitting where a new textile is created from scratch, or relying on textiles that exhibit rare properties. Oristitch hinges tuck away material between faces when heat-actuated through the use of a shrinking thread, similar to 4D Embroidery (Nabila et al., 2023a). The specific design of OriStitch hinges allows them to fully close and lock in place. By laying out hinges on a 2D sheet and closing them, OriStitch allows converting flat-fabricated sheets into 3D structure. We develop an Oristitch software tool based on Origamizer (Tachi, 2010) to automatically convert 3D meshes (OBJ) to stitch patterns, and compute an efficient division into hoops for fabrication. Exported files can directly be sent to laser cutters and embroidery machines. The resulting workflow allows creating 3D structures such as the cap shown in Figure 1 as well as other custom textiles like the vase cover and handbag shown in Figure 17. OriStitch integrates well with other textile manipulation techniques, such as quilting, embroidery, or sewing, allowing it to fit in with existing production workflows without reducing the expressiveness of design.
@@
Through technical validation, we find that OriStitch hinges fabricate and fold reliably using a range of materials (suede leather, cork, Neoprene, and felt). Our software tool successfully converted 26/28 models found in related papers (Guseinov et al., 2017; Narumi et al., 2023; Panetta et al., 2021; Jourdan et al., 2023) to machine instructions. With this contribution, we take a step to make the workflow of machine embroidery self-folding structures more accessible. The improved software support may allow fashion designers and discretionary users to design more elaborate patterns and advance the complexity of what they can make.
@@ 2. Contributions, Benefits, and Limitations
This paper makes three contributions. (1) We present a workflow, OriStitch, to create self-folding 3D geometry from flat fabrics. (2) We present a design of stitch patterns to achieve fully closing hinges under heat treatment. And finally, (3) we present a software tool to convert 3D meshes to fabrication instructions as well as an algorithm to efficiently compute hoop layouts. Combined, these contributions allow users to create a wide variety of shapes in a mostly automatic workflow.
Inherent to the use of developable surfaces, this process is limited to 3D objects that are topologically disks. We demonstrate that with the cutting of the 3D shape and stitching together separate sheets, more complex topologies (such as topological spheres and tori) are attainable. Furthermore, to be able to place hinges, objects are subject to a resolution of edge lengths of a minimum of . Our workflow uses high temperatures (¿350∘F), as well as exposure to water; therefore, the used fabric has to be robust to these conditions.
@@ 3.1. Textile Folding
Manual Processes. Manual techniques, such as hand pleating, allow folding a range of materials. During hand pleating, fabric is placed between two halves of a cardboard mold, folded, and afterwards it is exposed to high temperatures. Another technique is hand-sewing continuous threads into textiles and pulling them together. Despite providing greater design freedom, these processes are labor-intensive and time-consuming, limiting their practicality.
@@ 3.3. Techniques for Self-Folding
A wide variety of fabrication processes have recently been explored to achieve complex 3D shapes through 2D configurations. One of the most common self-folding techniques is to print thermoplastic composites using Fused Deposition Modeling (FDM) 3D Printers, also known as 4D printing (An et al., 2018). Within textiles, similar methods create self-folding structures by printing on pre-stretched sheets in a process called textile additive manufacturing (Brenken et al., 2018). Keefe et al. (2022) presents a great overview of these techniques, highlighting embroidery as one such process.
@@ 4.1. Stitch pattern design
Figure 3 shows the elementary hinge and its actuation principle. The hinges fold along the valley fold on the reflecting axis, so that the space between faces is completely tucked behind the surface. As a result, the decomposed faces will be brought together to form the target 3D geometry. To accomplish this, the active thread is made of a special polyester (PE) thread (Chizimi)222https://thesewingplace.com/chizimi-shrinking-thread-300m which shrinks by 30% of its original length when exposed to heat (350∘F). The two ends of the heat-shrinking thread are fixed on a pair of adjacent faces by so-called lock stitches. So that under heat treatment, as the active thread shrinks, the two faces will join. Each edge contains two parallel shrinking threads to uniformly close hinges. The so-called channel stitches guide the shrinking thread on the faces. The fold stitches are zigzag patterns with non-shrinking thread on top and shrinking threads on bottom (aka bobbin) for mountain folds, and vice-versa for valley folds. As the thread shrinks, the edge pre-creases in the predefined orientation, making it easier for the shrinking thread to close the hinge. Finally, the orange, water-soluble thread holds the active threads in place during fabrication.
@@ 4.1.1. Variation of hinge types
Between two faces, Oritstich inserts hinges, which tuck away excess material and ensure the faces meet up in the final 3D design. As a result the design of each hinge is dependent on the space between faces. The active thread shrinks by up to 30%, so widening the hinge means more of the active thread has to be stitched to each face. To accommodate for this, OriStitch has four different hinge types displayed in Figure 4. (b) The basic hinge (type 1) places the threads as laid out in Figure 3. (c-d) to accommodate for wider hinges, Oristitch has hinges of (type 2) which wrap the shrinking thread around both faces to share lock stitches, and (type 3) wraps two paths of shrinking thread next to each other all the way around each of the faces while inserting an additional pair of hinges to produce a harmonica fold and a shrinking thread that spans all the way around the other hinge. (a) Given the optimized layout of faces, there are faces that are so close that they do not allow for hinges to be inserted; here, Oristitch simply creases the edge between the faces (type 0).
@@
As illustrated in Figure 5a, each face has up to three neighboring faces, so they contain the shrinking patterns for up to three hinges. The fold stitches bend along the edges. This avoids intersections between the active threads, allowing them to share channel stitches and lock stitches. Besides simplifying the layout, sharing channel, and lock stitches speeds up fabrication.
4.1.2. Hinge type placement algorithm
The placement of each hinge is determined by a two-stage classification algorithm. This method evaluates the absolute width of a hinge and its width relative to the available folding space, which is a robust indicator of whether the material can physically accommodate the fold.
First, each side of a hinge (a and c) is independently assigned a Gap Type, , an integer score from 0 (No Gap) to 3 (wide). This score is a function of the hinge width and the available ”shrinking path” length . Figure 6 visually depicts these parameters in a prototypical hinge pair. The function prioritizes hinge sides that are already very narrow () and classifies no-gap type. For wider hinges, it calculates the critical ratio of width to available space, , and assigns a score based on predefined design thresholds (, , ):
Second, the hinge type is determined from the combination of the two sides’ gap types. By analyzing the sorted pair of scores, , the algorithm applies a set of rules that reflect practical manufacturing constraints. For example, a hinge is only classified as the type 0 if both sides are have no gap (), while a combination like is still acceptable and classified as type 1. This two-step process provides a nuanced and reliable assessment of each hinge’s feasibility.
Type 0 (No hinge): If
Type 1 (Narrow): If
Type 2 (Medium): If
Type 3 (Wide): If
Error: For all other combinations.
The way the Origamizer algorithm spaces out the faces from one another makes error cases of (0,2),(0,3), and (1,3) cases extremely rare. In our technical evaluation, we have encountered none of them. These would result in folding angles that our hinge design does not support. Users receive a warning when this happens.
4.1.3. constraints of the type 3 design
The active thread used in this paper shrinks by 30%, this constraints the design of hinges. For the long hinge design (type 3), the amount of material to fold into the harmonica fold depends on the total shrinkage. Here we lay out the way these constraints impact the design to get a sense for how to adjust to potential other threads. The basic premise is simple: we have an active thread, of length , and a hinge that tucks away material of height .
Let:
: the distance between anchor points on the two triangles
: the portion of the thread embedded inside the triangle
: the portion of the thread extending outside the triangle
: the total length of the thread used for this pair
To achieve the desired folding, the total shrinkage of the thread must equal . The thread we use has a shrinkage factor . From this geometric constraint:
Therefore:
We enforce two fabrication constraints:
1. To prevent the thread from extending too far and intruding into the opposite triangle, we ensure:
2. To ensure the thread sufficiently spans the hinge region and supports folding behavior, we require:
Solving for from both sides, we get:
| (6) | From the lower bound: | |||
| (7) | From the upper bound: |
Therefore, the hinge height must satisfy:
This expression provides a valid design window for choosing , ensuring both geometric feasibility and fabrication compliance. In our implementation, we select the hinge height as the midpoint of the feasible interval to balance thread coverage and prevent interference with adjacent triangles. Specifically, we choose:
By grounding the hinge parameters in both geometric constraints and thread material behavior, our method ensures that each hinge reliably contracts to the target folding height after shrinking, preserving the intended mechanical behavior and foldability.
Our next question is, by using this method of selection, will this satisfy all cases? Or is there any restriction we have here? The restriction is that the folding hinge h can not be too small , otherwise folding performance will not be guaranteed, though a smaller will support a longer hinge. Given our hinge design in 9, we impose the constraint:
Multiplying both sides by 2720 yields:
This inequality defines the minimal combination of anchor width and inner thread length required to maintain a hinge height of at least 2 mm. For usage with other threads we recommend characterizing the precise shrinkage and adjusting the parameters of these constraints accordingly.
@@ 4.1.4. Needle points
OriStitch’ hinge design requires careful placement of needle points, the stitches where bobbin threads (from below) connect with the threads on top. As illustrated in Figure 5b, in conventional embroidery, designers prioritize the top side of the fabric for aesthetic reasons, they provide extra tension on the bobbin thread, pulling it through the sheet. OriStitch cannot afford that because the active threads need to smoothly adjust their length. OriStitch adds additional needle points to distribute the tension across the thread, as shown in Figure 5c. This is specifically important at corners, where the thread on top makes a turn and would tear out the bobbin threads. OriStitch adds additional needle points at all corners and has less on the straight paths, ensuring both reliable fabrication and actuation.
@@ 4.1.5. Resolution
OriStitch is subject to a minimum size for each face to allow sufficient space to place all stitches. The minimal size of faces to be converted to OriStitch hinges has an edge length of 8.4mm and is predominantly constrained by embroidery machine limits. Typical machines have a needle point safety margin (distance between points) of 0.3mm, as shown in Figure 7, this is the main constraint for the size of faces in OriStitch. The channel stitch in the center is reduced to the minimal size, each corner requires three needle points, so on both sides of the active thread path there are exactly three points. The active thread path contains up to 6 parallel threads, which, in order for the threads to stay flexible, requires 0.6mm of width. The final constraining factor is the locks stitch design, which again for safety margin reasons cannot be further reduced. The main conclusion is that the resolution really is machine-specific when it comes to the safety margins. That said, this assumes a dense fabric (leather, felt, or materials with tightly woven structures), if the fabric has a structure with wider gaps (typical for loosely woven or knitted structures), we recommend increasing the minimum edge length in the model past this resolution before converting to stitch patterns. The Oristitch editor warns users when their OBJ file contains triangles that have edges below this minimum resolution, in this case users will have to re-triangulate their file.
@@ 4.2.1. Materials
The core material we used in this paper is a special polyester (PE) thread (Chizimi) that shrinks around 30% under heat treatment333https://thesewingplace.com/chizimi-shrinking-thread-300m/. We use water-soluble material444https://www.fjx.co.jp/english/catalog/household/water.html to temporarily hold other threads in place before actuation. Most of our testing has been done with ”Aida cloth, 18 count” fabric555https://www.michaels.com/product/10357725, but we find in the technical evaluation that the hinges continue to work with a range of fabrics. That said, caution should be placed on the selection of fabrics. During the actuation process both exposure to wetness and high temperatures limit the range of effective materials. We also found in our evaluation that the materials need to have some stiffness to work reliably, otherwise the hinges fail to fold as expected. OriStitch comes with a calibration patch shown in Figure 10, which we recommend to use for calibrating to material properties as well as actuation temperature.
@@
Next to machine-embroidery, OriStitch’s workflow makes use of laser cutting. As most fabrics are relatively flexible, they require scoring or pre-creasing of the edges to create well-defined folds with OriStitch. To reduce the force required to close the hinges and to create well-defined edges, mountain and valley folds cut along most of the edges, with only the top, middle, and bottom sections connected as demonstrated in Figure 4, Oristitch cuts away excess material between vertices to make it easier to fold. While the cutting step is not strictly required for the pattern to fold (as the original layout as generated by Origamizer (Tachi, 2010) folds paper without cutting), it simplifies the 2D layout as there is no need for additional patterns to fold the material away (aka ”vertex tucking molecules”), which in turn reduces the forces required to close each hinge.
As shown in Figure 8, OriStitch adds little tabs in the cutout areas to keep the sheet together; this minimizes the risk of the fabric folding on top of itself during embroidery and it keeps the tension in the sheet more even. Within the holes, OriStitch cuts small squares that are used for registration in the multi-hooping process. (b) OriStitch embroiders the same squares on the stabilizer sheet before attaching the fabric to the hoop. When attaching the fabric, these rectangles help align both sheets. Alternatively to laser cutting, some embroidery machines use cutting needles which can be used to remove the excess material. OriStitch exports both the laser cutting plan and stitch patterns as vector graphics (SVG).